The length of the rectangle is one meter less than twice its width. What are the dimensions of the rectangle?
The perimeter is 40 meters

Answer by rfer(2688)

2w+2(2w-1)=40
2w+4w-2=40
6w=42
w=7 meters
2w-1=13 meters